OpenStudy (anonymous):

a cube has a surface area of 6x^2+120x+600 find the length of a side

OpenStudy (turingtest):

I assume x represents the length of one side. The surface area of a cube = 6x^2 where x is the length of one side, because it has 6 faces, each with area x^2, so 6x^2=6x^2+120x+600 120x+600=0 x=-5 hmmm... I guess I would ignore the - sign and just say 5 but something seems wrong here.

OpenStudy (turingtest):

what does the x represent if not one side of the cube?

OpenStudy (anonymous):

you will have to find x for it i think so try this 6x^2+120x+600=0 take 6 common x^2+20x+100=0 x^2+10x+10x+100=0 x(x+10)+10(x+10)=0 x=-10 length never negative so lenght 10

OpenStudy (turingtest):

why do you think you can set this equal to zero ali?

jhonyy9 (jhonyy9):

x2 +20x+100 =0 x_1,_2=(-20 +/- sqrt(400-400))/2 =(-20+/- 0)/2= -10 - so lenght =10